OpenStudy (vishweshshrimali5):

If a,b,c, d and p are positive such that a+b+c+d + p= 8, a^2 + b^2 + c^2 + d^2 +p^2 = 16, then the maximum value of p is

OpenStudy (vishweshshrimali5):

a,b,c,d, and p are not necessarily natural numbers ; they are positive real numbers.

OpenStudy (vishweshshrimali5):

@ganeshie8 @hartnn

OpenStudy (anonymous):

can they be zero ? and real not integer /?

OpenStudy (vishweshshrimali5):

They cannot be zero. They are real but not natural numbers.

OpenStudy (anonymous):

ummm

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

they are positive and not 0 so 4^2 = 16 which makes it 1 number so others are less than 4

Miracrown (miracrown):

So this is a constrained optimization problem. non-linear thinking...

OpenStudy (anonymous):

so 3^2 = 9,2^2 =4 and rest would be 1 1 1

OpenStudy (anonymous):

so maximum value would be 3 dont u think

OpenStudy (anonymous):

aa they can be R not N

OpenStudy (anonymous):

oh

OpenStudy (jtvatsim):

but at the very least 3 is a lower bound for p

Miracrown (miracrown):

I am going to make the following argument. I am not actually certain that it's strictly speaking rigorous, but I think it will solve the problem here. Let's find out anyway. The argument is that there's no difference in the equations between a, b, c, and d. They all appear with the same constraints, same signs, and same functionality in the equations. So I argue that you can aggregate them together

OpenStudy (rational):

Maximize : \(f(b,c,d,p) = p\) subject to :\([8-(b+c+d + p)]^2+ b^2 + c^2 + d^2 +p^2 -16 = 0\)

OpenStudy (anonymous):

well

OpenStudy (rational):

lagrange multipliers gives the fast answer

OpenStudy (rational):

\[\langle 2[8-(b+c+d+p)] +2b, 2[8-(b+c+d+p)] +2c, \\ 2[8-(b+c+d+p)] +2d, 2[8-(b+c+d+p)] +2p\rangle \\= \lambda \langle 0,0,0,1\rangle \]

OpenStudy (rational):

I see lots of symmetry in the problem - we may use it if we can figure out a way - but above method gives you answer directly ^

OpenStudy (anonymous):

I'd probably say use cauchy–schwarz inequality for a,b,c and d

OpenStudy (anonymous):

have you tried it @vishweshshrimali5 ?

Miracrown (miracrown):

We start with the two constraints given. We have positivity a,b,c,d,p must be in the positive reals (excluding zero). Furthermore a+b+c+d+p = 8 a^2+b^2+c^2+d^2+p^2 = 16 What is the maximum that we can make p? Given this system. From the first equal and the positivity constraint, we can see that p = 8 - a - b - c - d. Since a,b,c,d>0. It follows that p < 8. So this is just a sanity check on any answers we may get. Now the symmetry argument because it's obviously hard to handle the four variables here. (We cannot use simplex here because we have a non-linear constraint a^2+...+p^2 = 16) The symmetry argument says, well, a,b,c,d have the same sign and make the same contribution to the constraint. So let's reduce them to a single variable "x" multiplied by 4. If we do that, we get a 2 variable system that we can solve. p = 8 - 4x p^2 = 16 - 4x^2 substitute p into the second equation (8-4x)^2 = 16 - 4 x^2 Expand 64 -64x + 16x^2 = 16 - 4x^2 Bring everything to one side 20x^2 - 64x + 48 = 0. Divide out GCF of 5, apply quadratic formula x = 16 +- Sqrt[16^2-4*5*12] all over 10 This gives x = 2, x = 6/5 x = 2 is not what we want because it makes p = 0 which violates the positivity constraint So we take x = 6/5 This yields p = 16/4 from p = 8 - 4x We check consistency of our answer to make sure that we did not mess up and violate the non-linear constraint 4x^2+p^2 = 16. We didn't. Okay, we have found the answer.

Miracrown (miracrown):

Well, if you argue that "a" "b" "c" and "d" are essentially 'the same' in the equation. Namely "x" then a^2 = x^2, b^2=x^2, and so a^2+b^2+c^2+d^2 = 4x^2

OpenStudy (vishweshshrimali5):

@Miracrown , your answer was absolutely correct (the answer should be 16/5). Here is how I got the answer, \[\large{(\cfrac{a+b+c+d}{4})^2 \le \cfrac{a^2 + b^2 + c^2 + d^2}{4}}\]

OpenStudy (jtvatsim):

ah... perhaps...

Miracrown (miracrown):

So the thing is. If you make "a" and little greater and "b" a little smaller in the hopes that you improve you value of "p" What would that gain you? Nothing it's the same as smoothing out the change over a single variable "x"

OpenStudy (anonymous):

A Mathematica constrained optimization solution is attached.

OpenStudy (anonymous):

according to cauchy–schwarz inequality\[4(a^2+b^2+c^2+d^2) \ge (a+b+c+d)^2\]using 2 constraints\[4(16-p^2) \ge (8-p)^2\]it gives\[p(5p-16) \le 0\]and finally\[0 \le p \le \frac{16}{5}\]

OpenStudy (jtvatsim):

wow, that is crazy cool how many ways there were to solve this thing! :)

Miracrown (miracrown):

Because a, b, c, d don't have a different coefficient in the constraints If they had different coefficients, it would be an entirely different and much harder problem

OpenStudy (vishweshshrimali5):

Yeah @mukushla , that's what I used.

OpenStudy (vishweshshrimali5):

Great work @mukushla and @Miracrown

OpenStudy (anonymous):

probably Cauchy–Schwarz inequality is the best bet for the problem

OpenStudy (anonymous):

3.2 close enough *wears glasses*

OpenStudy (jtvatsim):

lol @aajugdar

OpenStudy (jtvatsim):

could someone briefly explain how Cauchy-Schwarz applies? I've seen it used with dot product for the most part, is that hidden in here somewhere?

Miracrown (miracrown):

The thing is. There's no general method to solve these. It's just cleverness.

Miracrown (miracrown):

If you have systems of linear constraints, then you can use simplex But as soon as you get non-linear... it becomes tricky

Miracrown (miracrown):

So, cauchy-schwarz is basically the triangle inequality Do you know it by this name? @jtvatsim

OpenStudy (jtvatsim):

sure, |A + B| <= |A| + |B| right?

Miracrown (miracrown):

Yes, for the "2D" case

OpenStudy (jtvatsim):

ah, so it generalizes then.

Miracrown (miracrown):

exactly

OpenStudy (jtvatsim):

got it, so |A+B+C+...|^2 <= (|A|+|B|+|C|+...)^2

Miracrown (miracrown):

yes

Miracrown (miracrown):

So if a higher dimensional space, it's a longer length to go to A,B,C individually, then in some sense the straight line path from A--> X

OpenStudy (rational):

lagrange is the general method right @Miracrown

OpenStudy (jtvatsim):

ok, that seems reasonable that our notion of distance carries over

Miracrown (miracrown):

@rational Ah, you mean, if you want to use Langrangian multipliers? Actually, I did not think about that, but, yes, you could use them here

Miracrown (miracrown):

@jtvatsim yes

Miracrown (miracrown):

So cauchy-schwarz also guarantees tightness because, so you know that the distance can get no right than 4(a^2+b^2+c^2+d^2)

Miracrown (miracrown):

That's important, otherwise, you could have some really high upper bound which does not given you a meaningful constraint on p

OpenStudy (jtvatsim):

ok, I think I understand everything except where the 4 came from, is that just coming from the fact that we have 4 variables?

OpenStudy (jtvatsim):

thanks for taking the time to explain this much of the details :)

Miracrown (miracrown):

No problemo! Well... basically, but I'll explain

Miracrown (miracrown):

(sum i=1 to n of a_i*b_i)^2 must be <= (sum i =1 to n of ai^2) * (sum i = 1 to n of bi^2) This is the cauchy schwarz inequality phrased in terms of sequences of real numbers so if we let b = sequence of 1s (four ones) and a = sequence of our four variables then you see the 4 naturally comes from adding 1+1+1+1 for bi^2 thus (a+b+c+d)^2 <= 4*(a^2+b^2+c^2+d^2) That's where the 4 comes from

OpenStudy (jtvatsim):

ok, I'm copying and pasting this for further review. :)

Miracrown (miracrown):

(Actually this is the context of the first discovery of the cauchy-schwarz inequality in terms of sequences of real numbers) This is it's original form, I believe. Now-a-days, we mostly use it when dealing with vector spaces and such

OpenStudy (jtvatsim):

right, I usually think of vectors right away

OpenStudy (jtvatsim):

ah, great I see based on the sums that that would naturally follow. thank you!

Miracrown (miracrown):

you're welcome! :)